A polarizer and analyzer are oriented so that the amount of transmitted light is maximum. Through what




possible angles should either be turned so that the intensity of transmitted light by the analyzer is reduced




to 75% and 25% of the maximum intensity

1
Expert's answer
2022-01-19T19:28:01-0500

If "I_0"

is an incident, then after passing through

the polarizer then Intensity becomes "\frac{I_0}{2}"

Now when the analyzer is rotated the transmitted

intensity reduces.

"I=\frac{I_0}{2}cos^2\theta"

​

If θ is the angle of rotation

"\u03b8=30^0"

"I=\frac{I_0}{2}×cos^2 30
=0.375I_0"